OpenStudy (vipulrakta):

a ball is droped vertically from hight H on the plane surface and permitted to bounce repeatedly along the vertical line .after every bounce its kinetic energy became the quarter of its k.e. before the bounce the ball will came rest after how much time?

OpenStudy (osprey):

guess ... initial KE = KE1 ie BEFORE first bounce KE2 = KE1/4 KE3 = KE2/4 = KE1/4/4 = KE1/16 KE4 = KE3/4 = ... KE1/4/4/4 = KE1/64 .. KE5 = .... KE1/4/4/4/4 = KE1/256 ?????? Two poss answers it never comes to rest 'cos it's asymptotic it comes to rest when the above sequence renders the number too small to be measured. ?????? ps OpenStudy has flashed that it is closing down end of jan next year.